ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

8. Analysis

1991 MS Classification: not given Osmatescu, Petre, Moldova Basis of subtle spaces and some applications In work we expose the basis of the new notion of “subtle spaces”. The elements of subtle spaces are points and their particles. Some applications of subtle spaces are presented.